# SNOOPI: Supercharged One-step Diffusion Distillation with Proper Guidance

### v2

@@ Abstract
Recent approaches have yielded promising results in distilling multi-step text-to-image diffusion models into one-step ones. The state-of-the-art efficient distillation technique, i.e., SwiftBrushv2 (SBv2), even surpasses the teacher model’s performance with limited resources. However, our study reveals its instability when handling different diffusion model backbones due to using a fixed guidance scale within the Variational Score Distillation (VSD) loss. Another weakness of the existing one-step diffusion models is the missing support for negative prompt guidance, which is crucial in practical image generation. This paper presents SNOOPI, a novel framework designed to address these limitations by enhancing the guidance in one-step diffusion models during both training and inference. First, we effectively enhance training stability through Proper Guidance - SwiftBrush (PG-SB), which employs a random-scale classifier-free guidance approach. By varying the guidance scale of both teacher models, we broaden their output distributions, resulting in a more robust VSD loss that enables SB to perform effectively across diverse backbones while maintaining competitive performance. Second, we propose a training-free method called Negative-Away Steer Attention (NASA), which integrates negative prompts into one-step diffusion models via cross-attention to suppress undesired elements in generated images. Our experimental results show that our proposed methods significantly improve baseline models across various metrics. Remarkably, we achieve an HPSv2 score of 31.08, setting a new state-of-the-art benchmark for one-step diffusion models.
@@ 1 Introduction
Diffusion models have recently become popular due to their capacity to produce high-quality, diverse outputs in image [34, 39, 2, 52, 4], 3D [35, 48, 23], audio [17, 10] and video [53, 50, 3, 11] synthesis. Unlike other generative models, such as Generative Adversarial Networks (GANs), diffusion models gradually refine their output through a series of steps, allowing them to achieve diverse yet impressively detailed and realistic outputs. However, this iterative refinement is both time-consuming and computationally demanding, which limits their practicality in real-world applications. As a result, there’s increasing interest in techniques to speed up diffusion models while maintaining output quality. Several strategies are being explored to enhance their efficiency, including optimizing model architectures, improving sampling methods, and using latent space representations instead of image space. Recently, a particularly promising approach has gained traction: timestep distillation. Researchers have found that by adding specific constraints while training another model to mimic a pre-trained multi-step teacher, they can reduce the required steps, creating a faster and more efficient student that retains much of the original quality with minimal adjustments.
Figure 1: Comparison of various methods, including SNOOPI and alternatives (SwiftBrush, DMD, YOSO, PG-SB), using the PixArt-$\alpha$ backbone. Each method generates an image with the prompt ”Vincent van Gogh” (green). PG-SB helps stabilize the training process of image-free distillation, addressing the typical instability in such training. SNOOPI, a combination of PG-SB and the NASA module, enables effective negative prompting (red) even for one and few-step models, allowing for the removal of specific unwanted features and providing enhanced control over image attributes.
Current approaches for accelerating text-to-image diffusion models fall into two main categories: image-dependent training and image-free training methods. Image-dependent methods, which rely on real images or synthetic ones generated by the teacher models, include adversarial-based approaches [41, 19, 42, 51] and reconstruction-based methods [28, 56, 12, 19]. While effective, these methods demand huge storage, high memory, extensive computational power and might suffer from mode collapse due to too strong image supervision. In contrast, image-free methods [33], which mainly use score distillation losses like Score Distillation Sampling (SDS) [36] or Variational Score Distillation (VSD) [48], avoid the need for image datasets, instead requiring minimal data, often just text prompts. Models such as SwiftBrushv2 (SBv2) [7], even when omitting image-regularization term, making it a fully image-free distillation approach, have achieved impressive quality, even outperforming their multi-step counterparts on certain benchmarks. Hence, in this paper, we focus on the image-free approach and start with SBv2 as the baseline.
However, our research identifies two key challenges in current one-step distillation approaches using VSD. First, we observe training instability across different architectures, primarily due to the reliance on a fixed guidance scale. Second, we identify a broader issue affecting all one-step generators: the lack of negative prompt guidance, which limits the ability to exclude specific elements from generated images—a feature commonly supported in multi-step models [15, 8, 20]. These issues limit the flexibility and control of one-step models, which reduces their practical effectiveness despite their faster generation times.
Prior research [54] has attempted to address the stability issues of the VSD framework, but these solutions come with their own limitations. Techniques like increased teacher update frequency [54] and adversarial loss terms [42, 19] have been explored to stabilize the training process. However, these approaches can add a significant computational burden, partly due to the additional proposed technique itself and partly due to the introduction of image supervision. Since we’re focusing on building an image-free distillation, given these limitations, there is a pressing need to conduct a deeper analysis of the VSD framework to understand its challenges and develop more efficient, stable solutions.
Alongside these stability limitations, as previously mentioned, one-step diffusion models currently lack support for negative prompting, a feature widely used in multi-step diffusion models to enhance image quality and filter out unwanted features. In multi-step models, negative prompting techniques are particularly useful for refining alignment with text prompts in text-to-2D generation, and they also help address issues like the Janus problem in text-to-3D generation [8, 20, 1] —where models produce mirrored or duplicated features like faces. One common way to integrate negative prompts into multi-step diffusion model sampling is using Classifier-free Guidance (CFG) [15]. Given two noise predictions from the positive prompt and negative prompt, CFG subtracts negative noise prediction from the positive noise prediction iteratively throughout the denoising process, inherently steering the final output away from the negative prompt. However, one-step models do not have the iterative refinement process required by CFG, making it difficult to apply negative prompting in these frameworks. This limitation reduces the flexibility and control of one-step models. To address this, a novel approach is needed to incorporate the negative prompt signal directly within the one-step generation process, enabling these models to effectively avoid unwanted features without iterative sampling. We address these challenges by introducing a robust framework called SNOOPI for the one-step model, illustrated in Fig. 1, that enhances stability and control during both training and inference. (1) We propose Proper Guidance - SwiftBrush (PG-SB), a dynamic guidance approach that varies the guidance scale of teacher models during training to stabilize VSD. This flexible CFG improves the adaptability of image-free distillation across diverse diffusion backbones without adding computational or memory overhead. (2) We introduce a method called Negative-Away Steer Attention (NASA) used in inference, which incorporates negative prompt guidance into one-step models by adjusting cross-attention layers, effectively reducing unwanted features in generated images. By steering attention within intermediate feature space, NASA offers image control capabilities that were previously achievable only in multi-step diffusion models.
Our contributions are threefold:
  • We propose Proper Guidance-SwiftBrush (PG-SB), applying a varied guidance scale to teacher model outputs during training, to enhance stability across different diffusion backbones without additional computational costs.
@@
  • We introduce Negative-Away Steer Attention (NASA), the first method to integrate negative prompt guidance into one-step diffusion models, effectively suppressing unwanted visual attributes.
  • Extensive experimental results on established benchmarks demonstrate that our method outperforms strong baselines, validating its stability and effectiveness.
@@ 2 Related Work
2.1 Text-to-image Diffusion Distillation
Recent advances in text-to-image diffusion models have focused on accelerating generation through several key methodological streams: conventional distillation, trajectory optimization, and distribution matching.
Conventional Distillation. Early work in model distillation explored two main directions. Progressive distillation methods [40, 32] iteratively train a sequence of student models, each halving the required sampling steps while preserving generation quality. While effective, this approach necessitates multiple training phases, increasing overall computational cost. In parallel, direct distillation techniques [28, 12] generate synthetic noise-image pairs using a teacher model to train students via reconstruction loss, offering a more straightforward training process but often struggling with quality retention. Recent work has augmented these methods with adversarial training [19, 42], achieving better image quality at the cost of increased training complexity.
Trajectory Optimization. Trajectory-straightening methods [25, 26] simplify the complex denoising trajectories into near-linear paths, enabling faster sampling without compromising quality. Meanwhile, consistency-based frameworks [46, 21, 44] use self-consistency properties to train models to align outputs across multiple points in the ODE-based denoising trajectory.
@@ 2.1 Text-to-image Diffusion Distillation
Distribution Matching. More recent work [33, 7, 55, 54] has explored distribution matching techniques. These methods directly align the output distributions of teacher and student models, particularly through Variational Score Distillation (VSD). These techniques can be broadly categorized into two approaches: image-dependent training and image-free training.
@@
Image-dependent methods, which rely on real images for adversarial loss [54, 41] or synthetic ones generated by the teacher models for reconstruction loss [55]. Despite their effectiveness in preserving image quality, these approaches face significant practical limitations: they require substantial storage capacity, consume high memory during training, demand extensive computational resources, and risk mode collapse due to intensive image supervision.
@@
In contrast, image-free approaches [33, 7] have demonstrated promising results by relying solely on prompt text for training, offering a significant reduction in resource requirements while maintaining competitive performance. However, image-free techniques often encounter training instability, particularly within VSD-based frameworks. Existing solutions to stabilize training [55] introduce significant memory and computational overhead, limiting their practical application. Our work addresses this gap by introducing a simple yet effective method that enhances training stability without additional computational burden, making image-free variational score distillation more accessible for real-world applications.
2.2 Sampling Guidance and Negative Prompting
@@
Classifier-free guidance (CFG) [15] and negative prompting are key techniques for content control in image generation, from text-to-2D to text-to-3D synthesis. In multi-step diffusion models, these methods enhance desired image features and suppress undesired ones, improving visual fidelity and reducing artifacts like the Janus problem in 3D tasks [8, 20, 1]. Score-based approaches like NFSD [20] employ negative prompting to steer generated samples toward real-image distributions, while SDS-Bridge [31] uses negative prompts to model source distributions more accurately. Despite their success, negative prompting has yet to be integrated into one-step or few-step diffusion models. To bridge this gap, we introduce Negative-Away Steer Attention (NASA), the first method to enable negative prompting in one-step and few-step diffusion models, enhancing control in faster generation settings.
3 Background
@@
Diffusion Models [16, 45, 38] produce high-quality images by progressively denoising inputs. This technique involves a forward process that gradually adds noise to data and a reverse process that reconstructs the original data by removing noise. While the original diffusion models act on image space [16, 45], LDM [38] processes on latent space produced by a pretrained VAE for more efficient computation. In particular, starting with a data point $\mathbf{x}_{0}$ sampled from an unknown distribution $q(\mathbf{x}_{0})$, the forward process gradually diffuses $\mathbf{x}_{0}$ into a standard Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ through $T$ consecutive timesteps, where $\mathbf{I}$ is the identity matrix. At each time-step $t$, a noisier version of $\mathbf{x}_{0}$, denoted as $\mathbf{x}_{t}$, is drawn from $q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\alpha_{t}\mathbf{x}_{0},\sigma^{2}_{t}\mathbf{I})$ using a standard Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$:
@@
Classifier-free Guidance (CFG) [15] is an inference technique designed to enhance the quality of generated images by blending the predictions from both a conditional and an unconditional model. At each sampling step, CFG adjusts the denoiser’s output using a control parameter $\kappa>1$, allowing for controlled guidance that aligns more closely with the desired conditions:
@@
Negative Prompts provide enhanced control by suppressing unwanted features in the generated content. Instead of producing an unconditional output, the model generates an output conditioned on the negative prompt $\mathbf{y}^{-}$, as follows:
@@
Variational Score Distillation (VSD) is a powerful framework that utilizes pretrained diffusion-based text-to-image models to enhance text-based generation across both 3D and 2D domains. Originally proposed for text-to-3D tasks [48], VSD leverages diffusion-based score matching to align Neural Radiance Fields (NeRFs) with text prompt, enabling the generation of intricate 3D objects [48]. This approach extends effectively to 2D image synthesis as well, where VSD enables rapid, high-quality text-to-image generation in a single step with models such as SwiftBrush and DMD [33, 7, 55, 54], sidestepping the computational burden of multi-step diffusion methods.
The central aim of VSD is to ensures that the renderings of a differentiable generator align with the probability density of plausible images as guided by the 2D diffusion model. To accomplish this, VSD employs a two-teacher approach that uses a fixed pretrained diffusion model $\boldsymbol{\epsilon}_{\psi}$ and an adaptive LoRA-based teacher $\boldsymbol{\epsilon}_{\phi}$. While training, the student model $\boldsymbol{f}_{\theta}$ produces 2D images $\hat{\mathbf{x}}_{0}=\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})$ using an input noise $\mathbf{z}\sim\mathcal{N}(0,I)$ and a text prompt $\mathbf{y}$. Noisy images $\hat{\mathbf{x}}_{t}=\alpha_{t}\hat{\mathbf{x}}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ is then fed into both teacher models. The LoRA teacher model $\boldsymbol{\epsilon}_{\phi}$ aligns with the student distribution by minimizing a denoising L2 loss on single-step samples. This arrangement supports robust and adaptive guidance suitable for a variety of generative architectures. The gradient of the learning objective with respect to the student model parameters $\theta$ is defined as:
$\begin{split}\nabla_{\theta}\mathcal{L}_{\text{VSD}}=\mathbb{E}_{t,\mathbf{y},\mathbf{z},\kappa=C}&\bigg{[}\,w(t)\left(\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa)\right.\\ &\left.-\hat{\boldsymbol{\epsilon}}_{\phi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa\right)\frac{\partial\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})}{\partial\theta}\bigg{]},\end{split}$ (6)
@@
4 Proper Guidance - SwiftBrush
Figure 2: FID 2K progression throughout training. The default baseline is SwiftBrush for both SDv1.5 and PixArt-$\alpha$. DMD2Trick indicates the Two Time-scale Update Rule technique [54], which increases the update frequency of the LoRA teacher to 5 times after each student update. These baselines demonstrate notable instability, leading to variability in image quality metrics such as FID. In contrast, our approach, which employs CFG with randomly selected values from a uniform distribution, results in a more stable training process and generates higher-quality samples.

### v3

# Supercharged One-step Text-to-Image Diffusion Models with Negative Prompts

@@ Abstract
The escalating demand for real-time image synthesis has driven significant advancements in one-step diffusion models, which inherently offer expedited generation speeds compared to traditional multi-step methods. However, this enhanced efficiency is frequently accompanied by a compromise in the controllability of image attributes. While negative prompting, typically implemented via classifier-free guidance (CFG), has proven effective for fine-grained control in multi-step models, its application to one-step generators remains largely unaddressed. Due to the lack of iterative refinement, as in multi-step diffusion, directly applying CFG to one-step generation leads to blending artifacts and diminished output quality. To fill this gap, we introduce Negative-Away Steer Attention (NASA), an efficient method that integrates negative prompts into one-step diffusion models. NASA operates within the intermediate representation space by leveraging cross-attention mechanisms to suppress undesired visual attributes. This strategy avoids the blending artifacts inherent in output-space guidance and achieves high efficiency, incurring only a minimal 1.89% increase in FLOPs compared to the computational doubling of CFG. Furthermore, NASA can be seamlessly integrated into existing timestep distillation frameworks, enhancing the student’s output quality. Experimental results demonstrate that NASA substantially improves controllability and output quality, achieving an HPSv2 score of 31.21, setting a new state-of-the-art benchmark for one-step diffusion models.
Figure 1: Negative prompting are a key technique for improving controllability and image quality in text-to-image diffusion models. However, classifier-free guidance (CFG) [19] only supports negative prompt integration in multi-step models due to its iterative nature. Our method, NASA, is the first to enable negative prompting in one-step models, allowing for the suppression of unwanted features and providing greater control over image attributes.
00footnotetext: Equal Contribution
$\dagger$ Qualcomm AI Research is an initiative of Qualcomm Technologies, Inc.
@@ 1 Introduction
Diffusion models have recently become popular due to their capacity to produce high-quality, diverse outputs in image [37, 41, 3, 7, 36], 3D [38, 46, 26], audio [21, 15] and video [50, 5] synthesis. Unlike other generative models, such as Generative Adversarial Networks (GANs) [16, 23], diffusion models gradually refine their output through a series of steps, allowing them to achieve diverse yet impressively detailed and realistic outputs. However, this iterative refinement is both time-consuming and computationally demanding, which limits their practicality in real-world applications. As a result, there is an increasing interest in techniques to speed up diffusion models while maintaining output quality. Recent distillation methods [34, 13, 52, 51, 12] compress the entire multi-step generation process into a single step, a breakthrough that enables real-time image synthesis. This speedup makes diffusion models practical for applications like content creation and interactive media, where their slow performance was previously a major drawback. Despite these impressive speed advancements in one-step generators, we have identified a significant limitation that affects all these models: the inability to support negative prompting [48, 4]. This feature, which allows users to explicitly exclude unwanted elements, is widely available in multi-step diffusion models but remains notably absent in their single-step alternatives.
The most widely adopted approach for negative prompting, classifier-free guidance (CFG) [19] works by generating two parallel noise predictions during each denoising step – one incorporating the positive prompt and another using the negative prompt. The system then subtracts the negative prediction from the positive one, effectively steering generation away from unwanted attributes. This process relies on the iterative nature of multi-step diffusion, where corrections can be applied and refined over dozens of steps.
One-step generators, however, collapse this entire process into a single forward pass, eliminating the iterative refinement cycles that make CFG possible. Without these multiple correction opportunities, there is no straightforward way to incorporate negative guidance into the generation pipeline. Consequently, users face an unappealing choice between the instant generation that might include unwanted elements or waiting longer for more precisely controlled results. This speed-versus-control dilemma represents the most significant practical barrier to widespread adoption of one-step generators in professional creative workflows, highlighting the urgent need for novel approaches that can bridge this gap.
To address this issue, we introduce a novel method called Negative-Away Steer Attention (NASA) used in inference, which incorporates negative prompt guidance directly into one-step diffusion models. NASA works by adjusting cross-attention layers, effectively reducing unwanted features in generated images. By steering attention within intermediate representation space, NASA offers image control capabilities that were previously achievable only in multi-step diffusion models.
With our newly developed capability, we have explored integrating negative prompting directly into one-step distillation frameworks, focusing on the SwiftBrush (SB) family of models [34, 13]. The SB approach is compelling for its unique, image-free distillation methodology. By matching teacher model probability distributions and applying sophisticated training techniques, SBv2 can outperform its own teacher on standard image synthesis benchmarks. Despite these technical achievements, SB models still fall short in human preference evaluations. The core challenge lies in the distillation process: without real image supervision, the system lacks clear signals about which generation patterns should be avoided. This is precisely where negative prompting and our NASA framework come in.
Committed to maintaining SB’s valuable image-free distillation approach while incorporating guidance sampling to align with human preference, we provide the critical missing component, a way to steer the model away from undesirable outputs during training. NASA enables something previously impossible: guidance sampling within a one-step generator. Through this integration, our model achieves an HPSv2 [49] score of 31.21, establishing a new state-of-the-art on human preference metrics for one-step models.
In summary, our contributions are twofold:
@@
  • We introduce Negative-Away Steer Attention (NASA), the first method to integrate negative prompt guidance into one-step and few-step diffusion models in both inference and distillation settings.
  • By integrating NASA into the SwiftBrush (SB) distillation framework, we achieve a new state-of-the-art for one-step diffusion models on human preference metrics. This marks the first successful integration of negative prompt guidance directly into one-step model training.
@@ 2 Related Work
2.1 Distribution Matching Distillation
@@
Recent work [34, 13, 52, 51] has explored text-to-image diffusion distillation through distribution matching techniques. These methods directly align the output distributions of teacher and student models, particularly through Variational Score Distillation (VSD). These techniques can be broadly categorized into two approaches: image-dependent training and image-free training.
@@
Image-dependent methods rely on real images for adversarial loss [51, 42] or synthetic ones generated by the teacher models for reconstruction loss [52]. Despite their effectiveness in preserving image quality, these approaches face significant practical limitations: they require substantial storage capacity, consume high memory during training, demand extensive computational resources, and risk mode collapse due to intensive image supervision.
@@
In contrast, image-free approaches [34, 13] have demonstrated promising results by relying solely on prompt text for training, significantly reducing resource requirements while maintaining competitive performance. However, image-free techniques often encounter training instability, particularly within VSD-based frameworks. Existing solutions to stabilize training [52] introduce significant memory and computational overhead, limiting their practical application. Our work addresses this gap by introducing a simple yet effective method that enhances training stability without additional computational burden, making image-free variational score distillation produce better models.
2.2 Sampling Guidance with Negative Prompts
@@
Classifier-free guidance (CFG) [19] and negative prompting [48, 4] are key techniques for content control during denoising. In multi-step diffusion models, these methods enhance desired image features and suppress undesired ones, improving visual fidelity and reducing artifacts [14, 24, 1]. Score-based approaches like NFSD [24] employ negative prompting to steer generated samples toward real-image distributions, while SDS-Bridge [33] uses negative prompts to model source distributions more accurately. Despite their success, negative prompting has yet to be integrated into one-step or few-step diffusion models. To bridge this gap, we introduce NASA, the first method to enable negative prompting in one-step and few-step diffusion models, enhancing control in faster generation settings.
3 Preliminaries
@@
Diffusion Models [20, 44, 40] produce high-quality images by progressively denoising inputs. This technique involves a forward process that gradually adds noise to the data and a reverse process that reconstructs the original data by removing noise. While the original diffusion models act on the image space [20, 44], LDM [40] processes on the latent space produced by a pretrained VAE for more efficient computation. In particular, starting with a data point $\mathbf{x}_{0}$ sampled from an unknown distribution $q(\mathbf{x}_{0})$, the forward process gradually diffuses $\mathbf{x}_{0}$ into a standard Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ through $T$ consecutive timesteps, where $\mathbf{I}$ is the identity matrix. At each time-step $t$, a noisier version of $\mathbf{x}_{0}$, denoted as $\mathbf{x}_{t}$, is drawn from $q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\alpha_{t}\mathbf{x}_{0},\sigma^{2}_{t}\mathbf{I})$ using a standard Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$:
@@
Classifier-free Guidance (CFG) [19] is an inference technique designed to enhance the quality of generated images by blending the predictions from both a conditional and an unconditional model. At each sampling step, CFG adjusts the denoiser’s output using a control parameter $\kappa>1$, allowing controlled guidance that aligns more closely with the desired conditions:
@@
Negative Prompting [48, 4] provides enhanced control by suppressing unwanted features in the generated content. Instead of producing an unconditional output, the model generates an output conditioned on the negative prompt $\mathbf{y}^{-}$, as follows:
@@
Variational Score Distillation (VSD) is a powerful framework that utilizes pretrained diffusion-based text-to-image models to enhance text-based generation across both 3D [46] and 2D [34, 13, 52, 51] domains. The central aim of VSD is to align the renderings of a differentiable generator with the probability density of plausible images as guided by the 2D diffusion model. To accomplish this, VSD employs a two-teacher approach that uses a fixed pretrained diffusion model $\boldsymbol{\epsilon}_{\psi}$ and an adaptive LoRA-based teacher $\boldsymbol{\epsilon}_{\pi}$. While training, the student model $\boldsymbol{f}_{\theta}$ produces 2D images $\hat{\mathbf{x}}_{0}=\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})$ using an input noise $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and a text prompt $\mathbf{y}$. Noisy images $\hat{\mathbf{x}}_{t}=\alpha_{t}\hat{\mathbf{x}}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ are then fed into both teacher models. The LoRA teacher model $\boldsymbol{\epsilon}_{\pi}$ aligns with the student distribution by minimizing a denoising L2 loss on single-step samples. This arrangement supports robust and adaptive guidance suitable for a variety of generative architectures. The gradient of the learning objective with respect to the student’s parameters $\theta$ is defined as:
$\begin{split}\nabla_{\theta}\mathcal{L}_{\text{VSD}}=\mathbb{E}_{t,\mathbf{y},\mathbf{z},\kappa=C}&\bigg[\,w(t)\left(\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa)\right.\\ &\left.-\hat{\boldsymbol{\epsilon}}_{\pi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa\right)\frac{\partial\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})}{\partial\theta}\bigg],\end{split}$ (6)
@@
NP Anime Photo CA Paintings Average
29.62 29.17 28.79 28.69 29.07
31.66 29.17 30.58 30.38 30.42
Table 1: The effect of negative prompts (NP) on the HPSv2 score of PixArt-$\alpha$ [9]. Using generic negative prompts to remove visual artifacts (e.g., “worst quality”, “blurry”, “bad anatomy”) leads to a significant improvement. Higher scores indicate better alignment with human preferences.
4 Observations